A square coil is placed in a magnetic field. The coil is composed of 450 turns of wire. As the coil is wound around, the magnetic field changes at a constant tempo from 0,2 T to 0,42 T in 6 x 10-2 seconds. The induced emf is – 15 V.

  1. Calculate the change in magnetic flux in the coil.
  2. Calculate the length of the side of the coil
1
Expert's answer
2021-09-08T08:13:59-0400

Faraday's law of induction can be expressed as:


"\epsilon=-N\dfrac{d\Phi_B}{dt}=-NA\dfrac{dB}{dt}=-Nl^2\dfrac{dB}{dt}"


where "\epsilon" is the induced emf, N is the number of turns for the coil, and since it is a squared coil we also have A=l2 (length of a side to the square), and dB/dt is the rate of change for the magnetic field on the coil.


For the first par, we define and substitute for "\frac{d\Phi_B}{dt}" :


"\dfrac{d\Phi_B}{dt}=-\dfrac{\epsilon}{N}=-(\dfrac{-15\,V}{450\,turns})=(\dfrac{1\,V}{30})=0.0\overline{3}\frac{Wb}{s}"


After that, we use again the definition for Faraday's law and we find the square length l:


"\epsilon=-Nl^2\dfrac{dB}{dt} \iff l= \sqrt{\dfrac{-\epsilon}{N\dfrac{dB}{dt}}}
\\ \implies l= \sqrt{\dfrac{-(-15\frac{Wb}{s})}{(450)\dfrac{(0.42-0.20)T}{6\times10^{-2}s}}}
\\ l=\sqrt{\dfrac{(15\,m^2)}{(450)\dfrac{(0.42-0.20)}{6\times10^{-2}}}}=0.0953\,m"


In conclusion, we find that the change in magnetic flux in the coil "\dfrac{d\Phi_B}{dt}=0.0\overline{3}\,\cfrac{Wb}{s}", and the length of the side of the coil "l=0.0953\,m=9.53\,{cm}".


Reference:

  • Sears, F. W., & Zemansky, M. W. (1973). University physics.